OpenStudy (anonymous):

Write a system of equations for the following situation. Sharon has some $1 bills and some $5 bills. She has 14 bills which total to a value of $30.

OpenStudy (kuoministers):

let $1 = x and $5 = y ok? so 1x + 5y = 30 now that is one equation

OpenStudy (anonymous):

right

OpenStudy (kuoministers):

any idea what the 2nd one is?

OpenStudy (anonymous):

14x+14y=30?

OpenStudy (kuoministers):

x + y = 14

OpenStudy (kuoministers):

now we use a method to find x and y? which do you like ^^ elimination? substitution?

OpenStudy (anonymous):

the 1st one

OpenStudy (kuoministers):

awsome (so much easier) x + 5y = 30 x + y = 14 we subtract x from x vertically. 4y = 16

OpenStudy (anonymous):

y=4

OpenStudy (kuoministers):

now i hope you can figure it out. once you find y sub y into the equation

OpenStudy (anonymous):

isnt that for the substitution method?

OpenStudy (kuoministers):

nope

OpenStudy (kuoministers):

we use the elimination then we sub y into the equation

OpenStudy (kuoministers):

if we did substitution we would make y the subject then get y THEN we would sub y into equation to get x

OpenStudy (anonymous):

which equation should i plug the 4 in for y?

OpenStudy (kuoministers):

any :)

OpenStudy (kuoministers):

x + y = 14 would just be easier though

OpenStudy (anonymous):

ok ill go for that. I got 10 for x

OpenStudy (kuoministers):

good so we can conclude x = 10 and y = 4

OpenStudy (anonymous):

srry i didn't get the 2nd equation when u asked :(

OpenStudy (kuoministers):

mhm what equation?

OpenStudy (kuoministers):

x + y = 14?

OpenStudy (anonymous):

yea

OpenStudy (kuoministers):

well there is some $1 and $5 bills yea... and they add to 14. we said that $1 is x and $5 is y. so x + y is equal to 14.

OpenStudy (anonymous):

ok

OpenStudy (kuoministers):

thank you!!